OpenStudy (anonymous):

Solve the given equation for x. ln x^4 - 2ln x = 2. x=?

OpenStudy (anonymous):

how about this ?\[\huge \ln x^4 - 2\ln x = 2 \\ \huge 4\ln x - 2\ln x = 2 \\ \huge 2 \ln x=2 \\ \huge \ln x=1 \\ \huge x=e\]

OpenStudy (anonymous):

i used log properties to write \[\huge \ln x^4 =4\ln x \]

OpenStudy (anonymous):

yeah that looks better :)

OpenStudy (anonymous):

:D